I do not understand these at all. Ex. Jake paid $100.70 including 6% tax for a color printer. How much did the printer itself cost?
Answer by ikdeep(189)

Total money jack paid = $100.70
tax paid = 6% of cost of color printer.
Let the cost of printer = $c
so we get the equation...
total money = cost of printer + tax paid
$100.70 = $c + 6% of $c..when % is converted into fraction,,we get
100.70 = c + 6/100 * c ,,on dividing 6 by 100 ,,we get
100.70 = c + 0.06 * c ...now adding (c + 0.06 * c ) give us
100.70 = 1.06 c
or 100.70 / 1.06 = c

or c = 95
Therefore the cost of color printer = $95
hope this will help you
Please feel free to revert back for any further queries.